OpenStudy (anonymous):

Write an equation of the line that passes through the given point and is (a) parallel to the given line and (b) perpendicular to the given line. (6,0), y= 3/4x -1/4

OpenStudy (lgbasallote):

PARALLEL: slope = 3/4 equation of the line = y = 3/4(x-6) 4y = 3x - 18 3x - 4y -18 = 0 PERPENDICULAR: slope = - 4/3 equation of the line: y = -4/3(x-6) 3y = -4x + 24 4x + 3y -24 = 0

OpenStudy (anonymous):

Parallel: \[y=\frac{3}{4}(x-6), y=\frac{3}{4}x-\frac{9}{2}\] Perpendicular: \[y=-\frac{4}{3}(x-6),y=-\frac{4}{3}x-8\]

OpenStudy (anonymous):

\[y=-\frac{4}{3}x+8\], not minus, sorry about that

OpenStudy (anonymous):

THANKS <3

OpenStudy (anonymous):

The Parallel is always the same slope and then the Perpendicular is always the reciprocal of the slope. :)